# CKD: Contrastive Knowledge Distillation from A Sample-wise Perspective

### v1

@@ III-B Contrastive Knowledge Distillation
Generally, traditional knowledge distillation methods highlight class-level contrastive information and capture intra-class and inter-class relationships. Differently, our method focuses on sample-level representations. To be specific, the proposed approach aligns a student logit $\mathbf{s}_{i}$ with its corresponding teacher logit $\mathbf{t}_{i}$ while pushing it apart from other negative logits $\mathbf{s}_{j}$. To achieve this goal, we endeavor to distill not only intra-sample numerical similarities but also inter-sample semantic structure. Fig. 2 presents the overall architecture of CKD. We can see that our method conducts intra-sample distillation and inter-sample distillation. Afterward, we describe these constraints in detail.
@@ III-B1 Intra-Sample Distillation
Class-wise methods usually minimize intra-class distances among samples for knowledge transfer. By adhering to these methods, the proposed approach maximizes similarities between teacher and student logits for each sample while leveraging sample-wise information. Here, we utilize the intra-sample loss, denoted as $\mathcal{L}_{\text{intra}}$, to optimize parameters of the student model $\mathcal{S}$, where $d(\cdot,\cdot)$ represents a distance metric. Hence, we derive:
@@
For $\boldsymbol{x}_{i}\in\mathcal{X}$, we obtain $\boldsymbol{t}_{i}=\mathcal{T}\left(\boldsymbol{x}_{i};\theta_{\mathcal{T}}\right)$ and $\boldsymbol{s}_{i}=\mathcal{S}\left(\boldsymbol{x}_{i};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{x}_{i}\sim\mathcal{X}$ denotes a data point $\boldsymbol{x}_{i}$ sampled from $\mathcal{X}$. The loss function $\mathcal{L}_{\text{intra}}$, defined in Eq. 2, forces the student logit $\boldsymbol{s}_{i}$ to approximate its analytical solution $\boldsymbol{t}_{i}$. Previous studies [59, 55] have shown that continuously minimizing the gap between teacher and student representations can lead to overfitting and reduced generalization to unseen data. This is evident in the significant discrepancy between training and validation losses, especially when relying solely on intra-sample similarities.
@@
Fig. 3 provides a detailed analysis and visualization of the intra-sample constraint on logits $\boldsymbol{t}_{i}$, $\boldsymbol{t}_{j}$, and $\boldsymbol{s}_{i}$. The error between the $i$-th teacher and student logits is represented as $\epsilon_{i}=\boldsymbol{t}_{i}-\boldsymbol{s}_{i}$, where $\boldsymbol{t}_{i}\neq\boldsymbol{t}_{j}$. The objective $\mathcal{L}_{\text{intra}}$ cannot guarantee $\forall I,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$. Therefore, there exist errors $\epsilon_{i}$, such that ${\exists}i,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$. This inequality suggests that the student logit $\boldsymbol{s}_{i}$ may be closer to a teacher logit $\boldsymbol{t}_{j}$ than its corresponding teacher logit $\boldsymbol{t}_{i}$.
Formally, we assume that there are hyperplanes $\mathcal{H}^{(k)},k\in\mathbb{N}$ related to $\boldsymbol{s}_{i}^{(k)}\in\mathcal{H}^{(k)},i\in\mathcal{D}$ with the same loss value, that is:
$\forall k,h\;\;\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(k)}\right)=\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(h)}\right).$ (3)
The hyperplane $\mathcal{H}^{(k)}$ is defined as:
$\mathcal{H}^{(k)}:\mathbf{w}^{(k)}\boldsymbol{s}^{(k)}_{i}+b^{(k)}=0,\;i\in\mathcal{D},$ (4)
where $\boldsymbol{w}\in\mathbb{R}^{c}$ is a normal vector and $b$ is a bias.
Consider a two-dimensional plane $y=x$ and a teacher logit $\left[0.5,0.5\right]$. Despite student logits $\left[0.4,0.4\right]$ and $\left[0.6,0.6\right]$ yielding equal losses under the mean square error metric, they convey different meanings. Specifically, the logit $\left[0.4,0.4\right]$ lies below the decision plane, while $\left[0.6,0.6\right]$ lies above it. This highlights potential overfitting in intra-sample distillation. To address this, additional constraints are necessary.
@@ III-B2 Inter-Sample Distillation
Knowledge distillation should guarantee numerical similarity and semantic consistency between teacher logits and corresponding student logits [70, 4]. However, intra-sample alignment only emphasizes numerical similarity to align target logits. We introduce another inter-sample semantic alignment to simultaneously separate the logits of different samples. The inter-sample constraint $\mathcal{L}_{\text{inter}}$ is defined as:
@@
where $\boldsymbol{x}_{i}\neq\boldsymbol{x}_{j}$. The objective $\mathcal{L}_{\text{inter}}$ is to make $\boldsymbol{s}_{i}$ far from both $\boldsymbol{s}_{j}$ and $\boldsymbol{t}_{j}$. As shown in Fig. 3(b), we can alleviate the misalignment of $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{j}$ with Eq. 5 by maximizing their distance simultaneously.
Generally, inter-sample pairs include $\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$. We consider intra-sample pairs and can obtain triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$. Fig. 4 visualizes these two triple categories. Previous knowledge distillation approaches [70, 4] use both triple types for training. However, unlike these methods [64, 21, 8] in which each triple element is trainable, the teacher model remains fixed, with its output $\boldsymbol{t}_{i}$ unchanged. It may be questionable to simply apply these two triples in knowledge distillation methods.
To address this issue, we only use triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ during training. Fig. 4 briefly illustrates the rationale behind this selection. In Fig. 4(a), $\boldsymbol{s}_{i}$ approaches $\boldsymbol{t}_{i}$ for intra-sample alignment while distancing itself from $\boldsymbol{s}_{j}$ for inter-sample alignment. Even with a gradient conflict when $\theta_{i}>\pi/2$, we can resolve it by adjusting $\boldsymbol{s}_{j}$. However, in Fig. 4(b), the conflict remains unresolved due to the fixed $\boldsymbol{t}_{j}$. Additionally, Fig. 4(c) depicts the worst-case scenario where the gradient conflict is at its maximum. From Fig. 4, we can conclude that inter-samples are crucial for model updates and the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may cause gradient conflicts. Thus, our method merely adopt the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, and the inter-sample objective is formed as:
@@
Finally, we combine Eq. 2 and Eq. 6 with a tradeoff parameter $\beta$ to obtain the following loss function:
@@
Eq. 7 is to maximize similarities between teacher and student logits for the same sample while minimizing correlations across different samples. By pulling positive pairs closer and pushing negative pairs apart, the intra-sample and inter-sample formulations enable the alignment of teacher and student logits, ensuring they share similar semantic information for similarity and dissimilarity.
Figure 5: Pseudo code of CKD in a Numpy-like style.
@@ III-B3 Contrastive Formulation
The proposed structure-preserving alignment shares a similar objective with contrastive learning [54, 79]. We seek to transform our objective into a sample-wise contrastive learning formulation. Specifically, $\mathcal{L}_{\text{KD}}$ can be expressed as:
@@
For simplicity, we utilize the distance metric $f(\boldsymbol{u},\boldsymbol{v})$ to assess the similarity between $\boldsymbol{u}$ and $\boldsymbol{v}$, and set $f(\boldsymbol{u},\boldsymbol{v})=-d(\boldsymbol{u},\boldsymbol{v})$. The loss function can be derived as:
$\mathcal{L}_{\text{KD}}$ $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-(f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-\beta\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right))\right]$ (9)
$=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\ln\left[{\frac{\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}}+\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}-1\right]$ (10)
@@
The objective can be viewed as a sample-wise contrastive loss. Here, the teacher logit $\boldsymbol{t}_{i}$ and its corresponding student logit $\boldsymbol{s}_{i}$ form a positive pair, while $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ form a negative pair. Next, we will explore the benefits of this contrastive formulation.
(a) Traditional class-wise contrastive
(b) Typical class-wise contrastive: CRD
(c) Proposed sample-wise contrastive
@@
As shown in Eq. 9, to calculate the expected value $f(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ for each $\boldsymbol{x}_{i}$, we sample negative pairs from the data space $\boldsymbol{x}_{j}$. Following vanilla KD [25], we introduce a temperature parameter $\tau$ to soften the probability distribution, satisfying Eq. 12. The parameter $\beta$ balances the importance of positive and negative pairs, where a larger $\beta$ is assigned when negative pairs are important; otherwise, a smaller value is preferable. Note that negative sampling can also achieve this trade-off. When the negative space is significant, we increase the sampling of negative pairs for more detailed space estimation and weight adjustment. Conversely, fewer negative samples are selected when they are less important. Contrastive learning enables efficient batch data processing in parallel. In our formulation, we construct batches of training data, each containing one positive sample and $B-1$ negative samples, where $B$ is the batch size. From Fig. 2, the number of negative samples depends on $B$. Therefore, by varying $B$, we can sample different numbers of negative samples, balancing intra-sample and inter-sample alignment.
Additionally, we substitute $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ for $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. First, our method primarily aims to align $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$, such that $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}$. This leads to $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}\Rightarrow f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. Then, $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ involves two updating variables but $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ only has one variable, simplifying the optimization of $\boldsymbol{s}_{j}$. Specifically, we treat $\boldsymbol{t}_{i}$ as an anchor logit and separately learn its positive logit $\boldsymbol{s}_{i}$ and negative logit $\boldsymbol{s}_{j}$. Most importantly, this setup allows our method to leverage negative samples through the N-pair loss [64] or InfoNCE loss [54]. Similar to CLIP [58], we sample a batch of $B$ (teacher, student) pairs with $B\times B$ (teacher, student) pairings for efficient training.
@@
Given triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and the loss function $\mathcal{L}_{i}$ that is defined as:
@@
The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ and $\nabla_{\boldsymbol{s}_{j}}\mathcal{L}_{i}$ are proportional to $g_{i}$ that is formed as:
@@
For class-wise contrastive learning, $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ belong to different classes and $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=\rho_{j}$. When we decrease $B$, The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ also decreases.
@@
From Eq. 16, we can obtain following equations:
@@
When $B$ decreases, we sample fewer negative samples, resulting in a larger $\rho_{j}$ and a smaller $g_{i}$. Consequently, the gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ decrease. ∎
From Theorem 2, we can find that when positive samples and negative samples exhibit weak correlations, the value of $f(\boldsymbol{t}_{i},\boldsymbol{s}_{j})$ is minimized, reducing the gradient of backpropagation $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$. This hampers model parameter updates. Therefore, class-wise methods [64, 21] heavily rely on large batch sizes to mitigate this issue and improve learning efficiency [84].
@@
TABLE I: Classification accuracy (%) on the CIFAR-100 validation set, where teacher and student networks share the same architecture type. $\mathrm{\Delta}$ denotes the improved performance compared to vanilla KD. All results are the average over five trials.
Methods Teacher ResNet56 ResNet110 ResNet32$\times$4 WRN-40-2 WRN-40-2 VGG13
72.34 74.31 79.42 75.61 75.61 74.64
Student ResNet20 ResNet32 ResNet8$\times$4 WRN-16-2 WRN-40-1 VGG8
69.06 71.14 72.50 73.26 71.98 70.36
Features FitNet[59] 69.21 71 06 73.50 73.58 72.24 71.02
AT[31] 70.55 72.31 73.44 74.08 72.77 71.43
RKD[55] 69.61 71.82 71.90 73.35 72.22 71.48
CRD[70] 71.16 73.48 75.51 75.48 74.14 73.94
OFD[24] 70.98 73.23 74.95 75.24 74.33 73.95
ReviewKD[7] 71.89 73.89 75.63 76.12 75.09 74.84
Logits Vanilla KD[25] 70.66 73.08 73.33 74.92 73.54 72.98
TAKD[51] 70.83 73.37 73.81 75.12 73.78 73.23
DKD[90] 71.97 74.11 76.32 76.24 74.81 74.68
CKD 72.12 74.48 76.76 76.28 75.14 74.98
$\mathrm{\Delta}$ +1.46 +1.40 +3.43 +1.36 +1.60 +2.00
@@
For sample-wise contrastive learning, $\exists j$, $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ are in the same class, such that $\exists j,\;f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, large batch sizes are not crucial.

### v2

@@ III-B Contrastive Knowledge Distillation
Traditional knowledge distillation methods emphasize class-level contrastive learning by leveraging intra-class and inter-class relationships. In contrast, our approach adopts sample-level distillation. To be specific, for each sample, we enforce alignment between the student logit $\mathbf{s}_{i}$ with its corresponding teacher logit $\mathbf{t}_{i}$, while simultaneously distancing $\mathbf{s}_{i}$ from negative student logits $\mathbf{s}_{j}\;(j\neq i)$. This dual mechanism allows our method to distill not only numerical similarities in individual samples but also semantic structural relationships across different samples. As illustrated in Fig. 2, our contrastive knowledge distillation framework integrates two complementary components. The intra-sample distillation directly aligns paired student-teacher logits for each sample, and the inter-sample distillation employs contrastive separation between the target logit $\mathbf{s}_{i}$ and irrelevant logits $\mathbf{s}_{j}$. In the following sections, we elaborate on the formulation of these constraints.
@@ III-B1 Intra-Sample Distillation
Unlike class-wise methods that minimize intra-class distances for knowledge transfer, our method explicitly leverages sample-wise information to enhance the similarity between teacher and student logits for each individual sample. Here, we introduce the intra-sample loss, denoted as $\mathcal{L}_{\text{intra}}$, to optimize the parameters of the student model $\mathcal{S}$ by minimizing the distance $d(\cdot,\cdot)$ between aligned teacher-student logit pairs. Formally, the loss is defined as,
@@
For each input sample $\boldsymbol{x}_{i}\in\mathcal{X}$, the teacher and student models generate logits $\boldsymbol{t}_{i}=\mathcal{T}\left(\boldsymbol{x}_{i};\theta_{\mathcal{T}}\right)$ and $\boldsymbol{s}_{i}=\mathcal{S}\left(\boldsymbol{x}_{i};\theta_{\mathcal{S}}\right)$, respectively. $\boldsymbol{x}_{i}\sim\mathcal{X}$ means that a data point $\boldsymbol{x}_{i}$ sampled from the data space $\mathcal{X}$. In Eq. 2, the intra-sample loss $\mathcal{L}_{\text{intra}}$ enforces alignment between the student logit $\boldsymbol{s}_{i}$ and its analytical solution $\boldsymbol{t}_{i}$. While such alignment improves local similarities, earlier research [61, 56] reveals that exclusive reliance on intra-sample similarity risks overfitting as excessive minimization of the teacher-student gap may degrade generalization to unseen data. A phenomenon particularly evident in scenarios dominated by sample-specific optimization.
@@
Figure 5: Pseudo code of CKD in a Numpy-like style.
Fig. 3 visualizes the limitations of the intra-sample constraint by analyzing teacher logits $\boldsymbol{t}_{i}$ and student logits $\boldsymbol{t}_{j}$, where $\boldsymbol{t}_{i}\neq\boldsymbol{t}_{j}$. We define the alignment error for the $i$-th sample as $\epsilon_{i}=\boldsymbol{t}_{i}-\boldsymbol{s}_{i}$. While $\mathcal{L}_{\text{intra}}$ minimizes $d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})$, it does not guarantee the inequality $\forall i,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$. Consequently, there may exist cases where $\epsilon_{i}$, such that ${\exists}i,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$. This inequality implies that the student logit $\boldsymbol{s}_{i}$ becomes closer to an incorrect teacher logit $\boldsymbol{t}_{j}$ than to its assigned target $\boldsymbol{t}_{i}$.
Formally, we consider a family of hyperplanes $\mathcal{H}^{(k)},k\in\mathbb{N}$ in the student logit space, where each $\boldsymbol{s}_{i}^{(k)}$ under $\mathcal{H}^{(k)}$ satisfies $\boldsymbol{s}_{i}^{(k)}\in\mathcal{H}^{(k)},i\in\mathcal{D}$. These hyperplanes share the same aggregated loss value across all samples as,
$\forall k,h\in\mathbb{N}\;\;\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(k)}\right)=\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(h)}\right).$ (3)
Each hyperplane $\mathcal{H}^{(k)}$ is parameterized as,
$\mathcal{H}^{(k)}:\boldsymbol{w}^{(k)}\boldsymbol{s}^{(k)}_{i}+b^{(k)}=0,\;i\in\mathcal{D},$ (4)
where $\boldsymbol{w}^{(k)}\in\mathbb{R}^{c}$ is the normal vector and $b$ is the bias term.
Consider a binary classification scenario with a decision boundary defined by $y=x$. Suppose a teacher logs $\mathbf{t}_{i}=\left[0.5,0.5\right]$ corresponds to a sample near this boundary. Two student logits $\left[0.4,0.4\right]$ and $\left[0.6,0.6\right]$ yield identical intra-sample losses under mean squared error. However, their geometric positions relative to the decision boundary differ critically. $\left[0.4,0.4\right]$ falls below the boundary, implying misclassification. $\left[0.6,0.6\right]$ lies above the boundary, preserving correct classification. These loss values expose a key limitation of intra-sample distillation that $\mathcal{L}_{intra}$ may preserve numerical similarities while degrading task-specific correctness. To mitigate such overfitting risks, we introduce supplementary constraints that enforce semantic consistency beyond intra-sample alignment.
@@ III-B2 Inter-Sample Distillation
Knowledge distillation should guarantee both numerical similarity and semantic consistency between teacher and student logits [72, 4]. However, conventional intra-sample distillation focuses on the former by directly aligning paired logits. To address this limitation, we introduce inter-sample semantic distillation, which explicitly enforces structural relationships across samples. The inter-sample constraint $\mathcal{L}_{\text{inter}}$ operates complementarily to $\mathcal{L}_{\text{intra}}$ through contrastive separation by pushing student logits of distinct samples apart in the embedding space. Formally, $\mathcal{L}_{\text{inter}}$ is defined as,
@@
The objective loss $\mathcal{L}_{\text{inter}}$ enforces dual separation that it pushes the student logit $\boldsymbol{s}_{i}$ away from other student logits $\boldsymbol{s}_{j}$ and their corresponding teacher logits $\boldsymbol{t}_{j}$. As shown in Fig. 3(b), this constraint resolves cross-sample misalignment by preventing $\boldsymbol{s}_{i}$ from collapsing towards irrelevant teacher logits $\boldsymbol{t}_{j}$ and inter-sample redundancy by ensuring discriminative separation between distinct student logits $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$. Eq. 5 achieves this through simultaneous distance maximization between $\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$.
Generally, for each student logit $\boldsymbol{s}_{i}$, we construct inter-sample pairs include student-student pairs $\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ or student-teacher pairs $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$. We combine the intra-sample pair $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i}\right)$ and obtain two types of triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$. Fig. 4 visualizes these two triple categories. While previous triple-based methods [72, 4] leverage both triple types by updating all elements [66, 21, 8], our framework differs as the teacher model remains fixed with its outputs $\boldsymbol{t}_{i}$ unchanged during distillation. This constraint introduces a critical challenge that the direct adoption of traditional triplet losses designed for fully trainable models may lead to suboptimal alignment when applied to static teacher outputs.
Our method exclusively adopts $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ for distillation. Fig. 4 briefly illustrates the underlying reasoning behind this selection. In Fig. 4(a), the student logit $\boldsymbol{s}_{i}$ approaches $\boldsymbol{t}_{i}$ for intra-sample alignment while distancing from $\boldsymbol{s}_{j}$ for inter-sample separation. While gradient conflicts may emerge when the angle $\theta_{i}>\pi/2$, these can be mitigated by adjusting the trainable $\boldsymbol{s}_{j}$ during optimization. However, in Fig. 4(b), the conflict remains unresolved due to the immutable nature of $\boldsymbol{t}_{j}$. In extreme casesFig. 4(c), opposing gradients from $\boldsymbol{t}_{i}$ and $\boldsymbol{t}_{j}$ destabilize training, as neither term can be adaptively regularized. From Fig. 4, we conclude that inter-sample interactions must involve trainable student logits to ensure stable optimization and that the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may induce gradient conflicts. Consequently, our method exclusively adopts $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, and therefore formulate the inter-sample objective as the following equation,
@@
Finally, we unify the intra-sample alignment constraint Eq. 2 and inter-sample separation constraint Eq. 6 through a weighted summation by using a balancing parameter $\beta$ as,
@@
The objective Eq. 7 achieves intra-sample maximization by enhancing the similarity between teacher and student logits for the same sample and inter-sample minimization by suppressing correlations between the logits of distinct samples. This is realized through contrastive learning with positive pairs pulled closer via $\mathcal{L}_{intra}$ and negative pairs pushed apart via $\mathcal{L}_{inter}$.
@@ III-B3 Contrastive Formulation
Notably, the proposed structure-preserving distillation framework exhibits an objective akin to contrastive learning [55, 81]. This connection motivates us to reformulate our objective into a sample-wise contrastive learning framework. Formally, $\mathcal{L}_{\text{KD}}$ can be expressed as,
@@
For simplicity, we employ the distance metric $f(\boldsymbol{u},\boldsymbol{v})$ to quantify the similarity between $\boldsymbol{u}$ and $\boldsymbol{v}$, defining $f(\boldsymbol{u},\boldsymbol{v})=-d(\boldsymbol{u},\boldsymbol{v})$. The loss function can be derived as,
$\mathcal{L}_{\text{KD}}$ $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-(f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-\beta\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right))\right].$ (9)
where higher values indicate greater similarity. Employing Eq. 9 directly for optimization may result in gradient conflicts or diminish training efficiency. To rewrite it in the format of contrastive loss, we incorporate the natural logarithm for derivation as,
$\mathcal{L}_{\text{KD}}$ $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\ln\left[{\frac{\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}}+\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}-1\right]$ (10)
@@
Here, the teacher logit $\boldsymbol{t}_{i}$ and its corresponding student logit $\boldsymbol{s}_{i}$ form a positive pair, while $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ form a negative pair. Next, we will explore the benefits of this contrastive formulation.
(a) Traditional class-wise contrast
(b) Typical class-wise contrast: CRD
(c) Proposed sample-wise contrast
@@
As shown in Eq. 9, to compute the expected value $f(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ for each $\boldsymbol{x}_{i}$, we sample negative pairs from the data space $\mathcal{X}$. Following vanilla KD [25], we introduce a temperature parameter $\tau$ to soften the logit distribution defined in Eq. 12. The parameter $\beta$ dynamically adjusts the trade-off between intra-sample alignment and inter-sample separation. We assign larger $\beta$ values when semantic separation through negative pairs is critical and smaller values otherwise. Note that this balance can also be achieved through adaptive negative sampling. For example, in scenarios with dominant negative spaces, we increase the number of sampled negative pairs to improve space estimation and weight adjustment. When negative pairs are less informative, we reduce their sampling frequency. Contrastive learning enables efficient parallel processing of batch data. In our formulation, we construct training batches of data samples, each containing one positive sample and $B-1$ negative samples, where $B$ denotes the batch size. As shown in Fig. 2, the number of negative samples depends on $B$. Therefore, adjusting $B$ thus provides explicit control over the intra-inter sample balance.
Additionally, we replace $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ with $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. This substitution is justified by two key considerations. First, our method primarily aligns $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$, such that $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}$. We have $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}\Rightarrow f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. Second, $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ involves two trainable variables, whereas $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ depends only on $\boldsymbol{s}_{j}$. This optimization simplifies gradient propagation through $\boldsymbol{s}_{j}$. Specifically, we treat $\boldsymbol{t}_{i}$ as an anchor logit, and its positive logit $\boldsymbol{s}_{i}$ and negative logit $\boldsymbol{s}_{j}$ serve as learnable positive and negative logits, respectively. This design naturally connects with the InfoNCE loss [55], enabling batch-efficient training through $B\times B$ (teacher, student) pairings per batch analogous to CLIP [59]. Here, batch size $B$ controls the number of negative samples, balancing alignment accuracy and computational cost.
@@
Given triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$, the loss function $\mathcal{L}_{i}$ is defined as,
@@
The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ and $\nabla_{\boldsymbol{s}_{j}}\mathcal{L}_{i}$ are proportional to $g_{i}$ that is formed as,
@@
For class-wise contrastive learning, $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ belong to different classes and $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=\rho_{j}$. When decreasing $B$, the gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ also decreases.
@@
From Eq. 16, we can obtain following equations,
@@
When the batch size $B$ decreases, fewer negative samples are sampled, leading to an increase in $\rho_{j}$ and a decrease in $g_{i}$. This, in turn, results in a reduction of gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$. ∎
As proven in Theorem 2, when positive and negative samples exhibit weak correlations, the similarity measure $f(\boldsymbol{t}_{i},\boldsymbol{s}_{j})$ approaches its minimum. This suppresses the backpropagation gradient magnitude $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$, leading to inefficient model parameter updates. This limitation forces class-wise contrastive methods [66, 21] to rely on large batch sizes for gradient stabilization by amplifying signal-to-noise ratios through abundant negative samples and for correlation enhancement by mitigating weak correlation scenarios via batch sampling. Such strategies, while effective, incur substantial computational costs in contrastive learning [86].
@@
For sample-wise contrastive learning, $\exists j$ where $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ share the same class label, such that $\exists j,\;f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, large batch sizes are not crucial.